Question 24. Bank queues
bing, Exploring, and Comparing Data 23. Pulse Rates Listed below are pulse rates (beats per minute) from samples of adult males and females (from Data Set 1 Body Data in Appendix B). Does there appear to be a difference Male: 86 72 64 72 72 54 66 56 80 72 64 64 96 58 66 Female: 64 84 82 70 74 86 90 88 90 90 94 68 90 82 80 24. Bank Queues Waiting times (in seconds) of customers at the Madison Savings Bank an recorded with two configurations: single customer line: individual customer lines. Single Line 390 396 402 408 426 438 444 462 462 462 Individual Lines 252 324 348 372 402 462 462 510 558 600 Large Data Sets from Appendix B. In Exercises 25-28, refer to the indicated data set in Appendix B. Use software or a calculator to find the range, variance, and standard deviation Express answers using appropriate units, such as minutes. 25. Tornadoes Use the F-scale measurements from the tornadoes listed in Data Set 22 Torna does in Appendix B. Be careful to account for missing data. 26. Earthquakes Use the magnitudes (Richter scale) of the 600 earthquakes listed in Da Set 21 Earthquakes in Appendix B. In 1989, the San Francisco Bay Area was struck with earthquake that measured 7.0 on the Richter scale. If we add that value of 7.0 to those listed the data set, do the measures of variation change much? 27. Body Temperatures Refer to Data Set 3 Body Temperature in Appendix B and use t body temperatures for 12:00 AM on day 2 -28. Births Use the birth weights (grams) of the 400 babies listed in Data Set 4 Births to make an observation about those numbe How does that observation affect the way that the results should be rounded? Estimating Standard Deviation with the Range Rule of Thumb. In Exercises 29-32 refer to the data in the indicated exercise. After finding the range of the data, use the rang rule of thumb to estimate the value of the standard deviation. Compare the result to the standard deviation computed with all of the data
Answer #1


> male=c(86,72,64,72,72,54,66,56,80,72,64,64,96,58)
> mean(male)
[1] 69.71429
> sd(male)
[1] 11.65813
> female=c(64,8482,70,74,86,90,88,90,90,94,68,90,82)
> mean(female)
[1] 728.3077
> sd(feamle)
Error in is.data.frame(x) : object 'feamle' not found
> sd(female)
[1] 2329.716
> singleline=c(390,396,402,408,426,438,444,462,462,462)
> mean(singleline)
[1] 429
> sd(singleline)
[1] 28.6007
> inline=c(252,
+ inline=c(252,324,348,372402462462,510,558,600)
+ mean(inline)
Error: unexpected symbol in:
"inline=c(252,324,348,372402462462,510,558,600)
mean"
> inline=c(252,324,348,372,402,462,462,510,558,600)
> mean(inline)
[1] 429
> sd(inline)
[1] 109.2978
>